OpenStudy (anonymous):

If there are 0.505 g of NaCl left in a beaker that originally contained 75.0 mL of saltwater, what must have been the concentration of the original solution?

OpenStudy (anonymous):

NaCl m = 0.505 g v = 75 mL c= ? v= m/c Rearrange this to find c. c = v / m